OpenStudy (anonymous):

simplify the expression6+8^3

OpenStudy (anonymous):

is that 8 to the power of 3?

OpenStudy (anonymous):

Probably\[6+8^3=6+512=518\]

OpenStudy (anonymous):

6+ (8x8x8) 6+512 518